If you look at stock prices over any year, you will find a high and low stock price for the year. Instead of a single benchmark PE ratio, we now have a high and low PE ratio for each year. We can use these ratios to calculate a high and a low stock price for the next year. Suppose we have the following information on a particular company:
  

Year 1 Year 2 Year 3 Year 4
High price $ 87.06 $ 98.34 $ 119.70 $ 131.53
Low price 69.28 81.90 81.58 108.56
EPS 6.51 8.93 8.59 10.18

  
Earnings are projected to grow at 8.5 percent over the next year.  

What are the high and low PE ratios for each year? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)
  

High PE Low PE
Year 1
Year 2
Year 3
Year 4

  
What are the average high and low PE ratios over this period? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

High PE Low PE
Average

  
What is the high target stock price over the next year? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
  
High target price            $
  
What is the low target stock price over the next year? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
  
Low target price            $
Answer #1

a)

Year 1 Year 2 Year 3 Year 4
High price 87.06 98.34 119.7 131.53
Low price 69.28 81.9 81.58 108.56
EPS 6.51 8.93 8.59 10.18
High PE 13.37 11.01 13.93 12.92
Low PE 10.64 9.17 9.50 10.66

b)

Year 1 Year 2 Year 3 Year 4 Average
High PE 13.37 11.01 13.93 12.92 12.81
Low PE 10.64 9.17 9.50 10.66 9.99

c)

EPS 1 = EPS 0 (1+ g)

          = 10.18 (1 + 0.085)

          = 11.045

High target stock price over the next year = Average High PE * EPS 1

                                                                   = 12.81 * 11.045

                                                                   = 141.49

d) Low target stock price over the next year = Average Low PE * EPS 1

                                                                   = 9.99 * 11.045

                                                                   = 110. 34